Suppose consumers face the utility function:
U (Xi, Yt) = XiYt; for i = 1, 2, ... ... ... ..., M

Firms face the cost function:
Cj(qj) = q2j + F

(a) Solve for equilibrium in perfect competition.
(b) Show the effect of change in income on equilibrium quantity.
1
Expert's answer
2021-08-02T11:06:14-0400

a)

"U=xy\\M=p_xx+p_yy\\mu_x=y\\mu_y=x"

"\frac {mu_x}{p_x}=\frac{mu_y}{p_y}=\frac {y}{p_x}=\frac{x}{p_y}\\y=\frac{p_x}{p_y}X or X=\frac{p_y}{p_x}y"